OpenStudy (anonymous):

Graph f(x)= x(x+5)(x-3). Then find the area between the graph and the x-axis between x=-5 and x=3. area=?

OpenStudy (anonymous):

I don't know where to start

OpenStudy (gorv):

integrate it between x=-5 to 3 as integration gives area

OpenStudy (gorv):

\[\int\limits_{-5}^{3} (x^3+2x^2-15x)*dx\]

OpenStudy (gorv):

\[\frac{ x^4 }{ 4 } +2*\frac{ x^3 }{ 3 }-15*\frac{ x^2 }{ 2 } \]

OpenStudy (gorv):

now put value of integration limit in this \[3^4/4 +2*3^3/3-15*3^2/2-(-5)^4/4-2*(-5)^3/3-15*(-5)^2/2\]

OpenStudy (gorv):

81/4+18-135/2-625/4+250/3-375/2

OpenStudy (gorv):

@helfo just remember integration alwys give area